The silicon monosulfide diatomic molecule (SiS(g)) has a bond
dissociation enthalpy (BDE) of 619 kJ mol-1 , an inter-nuclear distance of
193 pm and an electronic dipole moment ( el ) of 1.73 Debyes.
a) Calculate the partial ionic character for SiS(g).

b) Compare the molecular orbital descriptions for the bonding in SiS (g)
Si2 (g) and S2 (g).
Answer #1

Given, dipole moment (observed) = 1.73 Debyes = 1.73 X 3.33564 X 10−30 C·m = 5.77 X 10-30 C.m

We know that dipole moment = electronic charge X interatomic distance

Now if the bond is 100% polarized then, dipole moment (100%) = (1.6 X 10−19) C X (1.41 X 10−10) m = 2.26 X 10−29 C m

ionic character = dipole moment (observed) /dipole moment (100%) = (5.77 X 10-30) / (2.26 X 10−29) = 0.26

% ionic character = 0.26 X 100 = 26%

For question b), it is too long to describe here. You need to draw three molecular orbital diagrams and probably you will also need to calculate the bond orders.